A coil receives an electric power of 4500 W from a supplied voltage of 240 V. What is the resistance of the coil?

1
Expert's answer
2022-03-26T15:31:15-0400

The power of electrical current

"P=\frac{V^2}{R}"

Hence, the resistance of the coil

"R=\frac{V^2}{P}=\frac{240^2}{4500}=12.8\:\Omega"